A father is 3 times as old as his son. In 12 years time, he will be twice as old. How old is the father now?
Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52776)

.

Son     x

Father 3x


In 12 years


    3x+12 = 2*(x+12).


    3x+12 = 2x + 24

    3x-2x  = 24-12

    x = 12.


Answer.  The son is 12 years old.  The father is 3*12 = 36 years old.

Solved.


Answer by Boreal(15235)

son is x
father is 3x
x+12 and 3x+12 where 3x+12=2(x+12)
3x+12=2x+24
x=12 son
father is 36
in 12 years 24/48